Let V be a finite dimensional inner product space over the field C of complex

numbers. Suppose B = {x1, x2, . . . , xn} is an orthonormal basis of V .

Prove that for all x, y ∈ V ,

(x|y) = Xn

i=1

(x|xi)(y|xi).


1
Expert's answer
2021-02-01T05:43:36-0500

If "B=\{u_1,u_2,...,u_n\}" is a basis of an n -dimensional inner product space "V" , then

"x=x_1u_1+x_2u_2+,...+x_nu_n"

"x=y_1u_1+y_2u_2+,...+y_nu_n"

So:

"\langle x,y \rangle=\langle \displaystyle \sum _{i=1}^nx_iu_i,\displaystyle \sum _{j=1}^ny_ju_j \rangle"